OpenStudy (anonymous):

The table shows how the distance traveled by a sailboat is changing over time. Time (in hours) Distance Traveled (in miles) 5 35 10 70 15 105 20 ? 25 175 What value is missing from the table? time is on the left distance on the right. answers are A. 110 B. 120 C. 140 D. 130

OpenStudy (alexandervonhumboldt2):

5/35=1/7 10/70=1/7 15/105=1/7

OpenStudy (alexandervonhumboldt2):

20/x=1/7

OpenStudy (alexandervonhumboldt2):

20/x=1/7 1/x=1/7*1/20 1/x=1/140 x=140

OpenStudy (acxbox22):

you can see from the table that every 5 hours you are going 35 miles more from this info, we can just add 35 to 105 because 20 hours is 5 more hours than 15 hours and the boat travels 35 miles in 5 hours

OpenStudy (anonymous):

ok I see how that works. Thanks :D

OpenStudy (acxbox22):

np ;)